OpenStudy (anonymous):

State whether the slope of a line passing through (8, 8) and (6, -12) is positive, negative, zero or undefined

OpenStudy (alexandervonhumboldt2):

again, use slope formula: \[\frac{ y_1-y_2 }{ x_1-x_2 }\]

OpenStudy (anonymous):

8- 12/ 9-6 = 1.333..

OpenStudy (alexandervonhumboldt2):

wrong

OpenStudy (alexandervonhumboldt2):

\[\frac{ 8-(-12) }{ 8-6 }=\frac{ 8+12 }{ 2 }\]

OpenStudy (alexandervonhumboldt2):

you forgot about negative

OpenStudy (anonymous):

6.666

OpenStudy (anonymous):

ooooh okay wait

OpenStudy (anonymous):

10

OpenStudy (anonymous):

so positive

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (alexandervonhumboldt2):

yewah it is 10